OpenStudy (anonymous):

1 slice pizza .75, 1 burger $1.35; 40 more slices of pizza sold than burgers. Sales totaled $292.50 How many slices pizza sold?

OpenStudy (imstuck):

you need to solve this with a system of equations. One equation relates the number of pizzas and burgers to the total number sold, and the other equation relates the money value of pizza and burgers to the total amount made.

OpenStudy (imstuck):

The first part of this tells you that a piece of pizza costs .75 and that a burger costs 1.35. It also tells you that the sales of both together is 292.50. So if pizza is P and burgers is B, then the money made for P plus the money made for B = 292.50. A piece of pizza is .75, so the money made from pizzas is .75P where P is the number of pieces of pizza times .75 each, and the money made from burgers is represented by 1.35B, where B is the number of burgers times 1.35 per burger. Those have to add up to = 292.50. So your first equation relates the money made in this way: .75P + 1.35B = 292.50

OpenStudy (imstuck):

Now for the numbers of each sold. You don't know much about either one of them but that the number of pieces of pizza sold was 40 more than the number of burgers sold. So P = 40 + B. Thats your second equation.

OpenStudy (imstuck):

Here they both are: P = 40 + B .75P + 1.35B = 292.50

OpenStudy (imstuck):

Since your first equation tells you that P is 40 + B, you sub 40 + B in for P in the second equation to get it in terms of B only.

OpenStudy (imstuck):

P = 40 + B .75P + 1.35B = 292.50 turns into this: .75(40 + B) + 1.35B = 292.50

OpenStudy (imstuck):

Now "do the math" and solve that for B. Can you do that?